Two equal point charges of +3.1x 10-6 C are placed 0.190 m apart. What are the magnitude and direction of the force each charge exerts on the other?
1
Expert's answer
2020-09-28T08:05:39-0400

Since both charges are of equal signs, the force acts to repel them. The magnitude of this force is


"F=\frac{1}{4\pi\epsilon_0}\frac{q\cdot q}{r^2}=\\\space\\
=\frac{1}{4\pi8.85\cdot10^{-12}}\cdot\frac{3.1\cdot10^{-6}\cdot3.1\cdot10^{-6}}{0.190^2}=\\\space\\
=2.39\text{ N}."